# How to train your VAE

### v2

@@ Abstract
Variational Autoencoders (VAEs) have become a cornerstone in generative modeling and representation learning within machine learning. This paper explores a nuanced aspect of VAEs, focusing on interpreting the Kullback–Leibler (KL) Divergence, a critical component within the Evidence Lower Bound (ELBO) that governs the trade-off between reconstruction accuracy and regularization [1, 2, 3]. Meanwhile, the KL Divergence enforces alignment between latent variable distributions and a prior imposing a structure on the overall latent space but leaves individual variable distributions unconstrained. The proposed method redefines the ELBO with a mixture of Gaussians for the posterior probability, introduces a regularization term to prevent variance collapse, and employs a PatchGAN discriminator to enhance texture realism. Implementation details involve ResNetV2 architectures for both the Encoder and Decoder. The experiments demonstrate the ability to generate realistic faces, offering a promising solution for enhancing VAE-based generative models.
@@ 1 Introduction
The KL Divergence has a pivotal role in VAEs as it enforces alignment between the distribution of latent variables and a prior distribution, often assumed to be a Normal distribution with zero mean and unit covariance. This divergence term promotes the emergence of structured latent representations, a cornerstone of effective generative models. However, our investigation delves deeper investigate on the latent space structure. We note that the KL Divergence does not impose any constraints on the individual distributions of the latent variable associated with a datum (individual posterior); this has implications for the diversity of latent vector behaviors in VAEs. In light of these findings, our proposal reformulates the role of KL Divergence with implications for latent space modeling. Our proposal addresses a well-known challenge in VAEs: the posterior collapse phenomenon [10], which can lead to a loss of distinctiveness in generated data. Specifically, we demonstrate the successful application of our approach in generating realistic facial images while retaining crucial variations in features like hairstyles and specific facial attributes [11, 8]
The main contributions in this work are an ELBO that reduces the collapse of the posterior towards the anterior (observed as the generation of very similar, blurry images). Our ELBO can be seen as an opposite version of Beta-VAE that preferentially weights quality reconstructions and an regularization term that prevents collapse of individual posteriors to a Dirac delta. Additionally, we use a VAE-GAN type scheme that promotes realistic data. Finally, we employ a Residual type architecture for the encoder and decoder to increase the networks capabilities. The main contributions affect the training process, so our results must be compared directly with those of a standard VAE [1, 4] or Beta-VAE [7]. Qualitative comparisons of the presented results versus those reported in the classic VAE or Beta-VAE formulations show superior performance of our proposal.
@@
We organize this paper as follow. We present in Section 2 a brief review of the standard VAE model. Then, in section 3, we present a mixture of Gaussian’s for defining the posterior probability $q({\bf z}|{\bf x})$ in the VAE, which leads us to redefine the ELBO, and we present practical implementation steps. Thus, we offer experimental results that support our claims in section 4. By advancing our understanding of the regularization mechanism in VAEs and its practical implications, we contribute to the ongoing pursuit of enhancing generative modeling techniques for complex data distributions. Through this Work, we hope to shed light on the interplay between regularization and diversity in latent spatial representations, contributing to the advancement of generative models in machine learning and artificial intelligence.
@@ 2 Background
The goal is to learn the generative model of $\mathcal{D}$} a dataset with $N\geq 1$ independent and identically distributed datapoints: ${\mathcal{D}}=\{{\bf x}^{(i)}\}_{i=1}^{N}$ with ${\bf x}^{(i)}\in\mathbb{R}^{m}$; assuming a set of latent variables, one for each datapoint: ${\mathcal{Z}}=\{{\bf z}^{(i)}\}_{i=1}^{N}$ with ${\bf z}^{(i)}\in\mathbb{R}^{d}$, and $d<m$. In this way, we have $({\bf x},{\bf z})\sim p({\bf x},{\bf z})=p({\bf x}|{\bf z})p({\bf z})$, which suggests a two-step generative model:
@@
$\mathrm{ELBO}(\phi,\theta)=$ $\beta\,\mathrm{KL}\bigl(q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr)$ (3)
$-\,\mathbb{E}_{q_{\lambda}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]$ (4)
where the parameter $\beta$ weights the relative contribution of each term to the total cost. The standard VAE formulation uses $\beta=1$, meanwhile $\beta$-VAE propose to use $\beta_{1}>1$ and $\beta_{2}=1$ to improve interpretable factorised latent representations; i.e., to learn disentangled latents [11]. The second term (4) is the negative log-likelihood that enforces the generated data, $\hat{\bf x}^{(i)}$, to be likely to belong to the training dataset. The first term (3) is Kullback-Leibler divergence that enforces the approximated posterior $q_{\phi}$ to be similar to a chosen prior $p_{\lambda}$; i.e. ${\bf z}\sim p_{\lambda}({\bf z})$ and $\lambda$ are the parameters of the prior. The most common approach is to assume the prior distribution $p_{\lambda}({\bf z})$ a multivariate Gaussian with a mean of zero and an identity covariance matrix:
@@
Hence, the one can factorize the posterior as [3]:
@@
This factorization takes advantage of the fact that the product of Gaussians is a Gaussian. The second row is the product of the individual posteriors (associated with each data point ${\bf x}^{i}$), and the third row makes explicit the marginals (of each latent dimension ${\bf x}_{j}$). Fig. 2 illustrates the Gaussian mixture model for the posterior. The dashed line represents the posterior $q_{\phi}$; i.e., the product of the individual Gaussian $q_{\phi}^{(i)}$ associated with each data point (in color). Note that while $q_{\phi}$ is enforced to be similar to the prior $p_{\lambda}$: a Gaussian with mean equal to zero and unitary variance. Since the KL between two univariate Gaussian is given by
@@
$\mathrm{KL}\bigl(q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr)=\sum_{j=1}^{d}\mathrm{KL}\bigl(\mathcal{N}(\bar{\mu}_{j},\bar{\sigma}_{j})\|\mathcal{N}(0,I)\bigr)\\ =\frac{1}{2}\sum_{j=1}^{d}\sum_{i:{\bf x}^{(i)}\in D}\left[(\sigma_{j}^{(i)})^{2}+(\mu^{(i)}_{j})^{2}-1-\log(\sigma_{j}^{(i)})^{2}\right];$ (10)
where $\bar{\mu}$ and $\bar{\sigma}=\mathrm{diag}(\sigma_{j}^{2})$ are the mean vector and covariance diagonal matrix of the approximated posterior (7), respectively. The ELBO minimization is efficiently achieved by using stochastic gradient descent by using samples of the training set (mini-batches): $\mathcal{M}\subset\mathcal{D}$.
@@ 3.1 The posterior collapse problem and over–smoothed reconstructions
VAE often suffer of the posterior collapse problem [12] (latent variable collapse or KL vanishing), such a problem occurs when $\mathrm{KL}\bigl(q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr)\rightarrow 0$. The posterior collapse to the prior results in non–identifiable latent variables: the likelihood function is independent of the latent representation producing a smoothness of the reconstructed images. To alleviate such problem there has been suggested many possible solutions, see Ref. [13] and references therein. Despite of many efforts to alleviate the posterior collapse, one can yet observe over-smoothed predictions [7, 12, 13]. It seems that only hierarchical computational expensive VAE architectures are capable of generating realistic textured data [14, 5].
To reduce the blurring, a naive strategy could consists in to reduce the parameter $\beta<1$ in (3) and passing the weight to the log-likelihood. This present another problem that we will analyze following. Note that the covariance $\bar{\sigma}$ of the posterior $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ is a diagonal (non–correlated) matrix the product in (8) can be decomposed in products of univariate Gaussians, one product for each dimension $j$. Thus, the jj–th product results in
@@
and $s_{j}$ is a scale factor which exact form can found in [15]. If in order to obtain reconstructions that are more faithful to the data, we weight more log–likelihood term into the ELBO then the individual posterior $q_{\phi}^{(i)}$ variance is reduced so the variations of the predictions are limited such that the VAE conducts close to a simple Autoencoder. In such case, Equations (11)–(13) make evident a problem that arise when the individual posterior $q_{\phi}^{(i)}$ approaches to a Dirac delta: $\sigma_{j}^{(i)}\approx 0$. Thus, the jj–th entry of the posterior $q_{\phi}$ [see (7)] approaches to a Dirac Delta: $\bar{\sigma}_{j}\approx 0$. As result, the $z_{j}$ latent variable is meaningless. Therefore it is to be expected that a significantly large $\beta$ value results in more blurred reconstruction.
@@ 3.2 The posterior as a mixture of Gaussians
To introduce our proposal, first, we note that the latent variable ${\bf z}^{(i)}$, associated with the datapoint ${\bf x}^{(i)}$, is also a stochastic variable with its particular distribution, the individual posterior:
@@
where $q^{(i)}_{\phi}({\bf z})=\mathcal{N}(\mu^{(i)},\sigma^{(i)})$ with $\sigma^{(i)}\succ 0$. Such individual posteriors need to be combined to define the posterior. A direct possibility is the factorization (7) but others factorization are possible. In this Work, instead of the product (7), we propose the mixture model:
@@
with $\alpha_{i}=1/N$, which implies that all data points are equally relevant. Hence, the mean of the mixture is given by:
@@
Thus, we can estimate the mean and variance of the global posterior with a Stochastic Gradient strategy using data batches $\cal D$}:
$\bar{\mu}_{j}$ $\approx\frac{1}{N}\sum_{i:{\mathbf{x}}_{i}\in{\cal D}}\mu^{(i)}_{j}$, (17)
$\bar{\sigma}^{2}_{j}$ $\approx\frac{1}{N}\sum_{i:{\mathbf{x}}_{i}\in{\cal D}}\left({\mathbf{z}}^{(i)}_{j}-\bar{\mu}_{j}\right)^{2}.$ (18)
@@ 3.3 Loss and training strategy
In this section we present our proposal of loss. We can note that the classic $\mathrm{KL}$ formulation [1] encourages the individual means $\mu^{(i)}$’s to be zero, contributing to the so-called posterior collapse effect: the ${\mathbf{z}}^{(i)}$ are non–identifiable. Differently, in our proposal, we encourage the mean of these $\mu_{i}$’s to approach zero:
$\mathrm{KL}_{G}\bigl(q_{\phi}(\mathbf{z}|\mathbf{x})\,\|\,p_{\lambda}(\mathbf{z})\bigr)=\mathrm{KL}\bigl(\mathcal{N}(\bar{\mu},I\bar{\sigma}^{2})\,\|\,\mathcal{N}(0,I)\bigr)\\ =\frac{1}{2}\sum_{j=1}^{d}\left[\bar{\sigma}_{j}^{2}+\bar{\mu}_{j}^{2}-1-\log\bar{\sigma}_{j}^{2}\right];$ (19)
where $\bar{\mu}$ and $\bar{\sigma}^{2}$ are given in (17) and (18), respectively.
Furthermore, one behavior we aim to prevent is the collapse of the variances of individual posteriors $\sigma^{2}_{i}$ to zero. That eliminates the stochastic behavior of the Encoder, causing the VAE to behave like an AE. In such a scenario, the VAE would be limited to generalizing: to generate data that were not part of its training data [16].Thus, we introduce a regularization term that penalizes individual variances that become close to zero: avoiding such individual posteriors to become Dirac deltas. This regularization term is:
@@
Note that KL (20) does not penalizes the individual means.
It is well known that metrics based on pixel-to-pixel comparisons are not the most appropriate to evaluate the quality of data produced by generating models. We are willing to tolerate significant reconstruction errors as long as the reconstruction looks realistic. In the example we use, this means that we prefer to generate images where the hairstyle seems realistic, even if the positions of the curls in the reconstruction do not match the input image, rather than over-smoothing regions of the hair (as is common in a typical VAE). GAN-type models have been relatively successful in generating realistic images [17], despite suffering from the problem of modes collapse. Similar to VAE-GAN that trains the VAE based on a GAN-type scheme [18], we also use a GAN training scheme. In particular, we use a PatchGAN strategy [19, 20, 21], which instead of giving a reality measurement rating for the entire image, generates a two-dimensional array where each element evaluates a support region in the generated image or real.
Therefore the likelihood term includes two metrics, a pixel to pixel that enforce the restoration to be similar to the input data and a general evaluation grade of the image realism. For the first term we choose the L1 norm known for being robust to errors in the exact reconstruction of structures’ positions in high-frequency texture areas. This is equivalent to said that the error follows a Laplace distribution. For the second term we use a Convolutional Neural Network (Discriminator) that given the restored image $\hat{\mathbf{x}}$, of size (H,W), produces an output $\mathbf{r}$}, of size $(h/8,w/8)$, where each entry reflects the realism degree of its support region. The discriminator is represented by the conditional ${\mathbf{r}}\sim p_{\omega}({\mathbf{r}}|{\mathbf{x}})$ where ${\mathbf{r}}\in[0,1]^{h/8\times w/8}$, and $\omega$ are the parameters of the neural network. Then the log–likelihood takes the form
[…]
@@
[…]
This loss encourages the generator to produce realistic predictions $\hat{\mathbf{x}}$. In order to train the discriminator we minimize the loss
[…]
Fig. 3: Residual-VAE components. (a) Residual encoder, (b) Gaussian sampler, and (c) residual decoder.
In all the conducted experiments we kept the weight parameters: $(\beta_{1},\beta_{2},\beta_{3},\beta_{4})=(1,0.5,1000,100)$. For computational efficiency purposes, we implement the train of both, the VAE (generator) at the discriminator, using stochastic descend strategy with batch size equal $b$. Note that as the batch size $b$ becomes larger, the estimation of ELBO terms [Eqs. (19), (20) and (21)] becomes more precise. Therefore, avoiding using tiny batches is important to ensure stable convergence of the stochastic gradient. Algorithm 1 summarizes training details.
@@
2:  ${\mathbf{z}}\sim q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ $\triangleright$ Encode
3:  $\hat{\mathbf{x}}\sim p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ $\triangleright$ Decode
[…]
@@
6:  Take a gradient descent step on $\nabla_{\omega}L(\omega)$.

### v3

@@ Abstract
Variational Autoencoders (VAEs) have become a cornerstone in generative modeling and representation learning within machine learning. This paper explores a nuanced aspect of VAEs, focusing on interpreting the Kullback–Leibler (KL) Divergence, a critical component within the Evidence Lower Bound (ELBO) that governs the trade-off between reconstruction accuracy and regularization. Meanwhile, the KL Divergence enforces alignment between latent variable distributions and a prior imposing a structure on the overall latent space but leaves individual variable distributions unconstrained. The proposed method redefines the ELBO with a mixture of Gaussians for the posterior probability, introduces a regularization term to prevent variance collapse, and employs a PatchGAN discriminator to enhance texture realism. Implementation details involve ResNetV2 architectures for both the Encoder and Decoder. The experiments demonstrate the ability to generate realistic faces, offering a promising solution for enhancing VAE-based generative models.
@@ 1 Introduction
The KL Divergence has a pivotal role in VAEs as it enforces alignment between the distribution of latent variables and a prior distribution, often assumed to be a Normal distribution with zero mean and unit covariance. This divergence term promotes the emergence of structured latent representations, a cornerstone of effective generative models. However, our research delves deeper into the latent spatial structure. We note that the KL Divergence does not impose any constraints on the individual distributions of the latent variable associated with a datum (individual posterior); this has implications for the diversity of latent vector behaviors in VAEs. In light of these findings, our proposal reformulates the role of KL Divergence with implications for latent space modeling. Our proposal addresses a well-known challenge in VAEs: the posterior collapse phenomenon [8], which can lead to a loss of distinctiveness in generated data. Specifically, we demonstrate the successful application of our approach in generating realistic facial images while retaining crucial variations in features like hairstyles and specific facial attributes [9, 6]
The main contribution in this work is an ELBO that reduces the collapse of the posterior towards the anterior (observed as the generation of very similar, blurry images). Our ELBO can be seen as an opposite version of Beta-VAE that preferentially weights quality reconstructions and a regularization term that prevents the collapse of individual posteriors to a Dirac delta. Additionally, we use a VAE-GAN type scheme that promotes realistic data. Finally, we employ a Residual type architecture for the encoder and decoder to increase the network capabilities. The main contributions affect the training process, so our results must be compared directly with those of a standard VAE [1, 2] or Beta-VAE [5]. Qualitative comparisons of the presented results versus those reported in the classic VAE or Beta-VAE formulations show the superior performance of our proposal.
@@
We organize this paper as follow. We present in Section 2 a brief review of the standard VAE model. Then, in section 3, we present a mixture of Gaussian’s for defining the posterior probability $q({\bf z}|{\bf x})$ in the VAE, which leads us to redefine the ELBO, and we present practical implementation steps. Thus, we offer experimental results that support our claims in section 4. By advancing our understanding of the regularization mechanism in VAEs and its practical implications, we contribute to the ongoing pursuit of enhancing generative modeling techniques for complex data distributions. Through this work, we hope to shed light on the interplay between regularization and diversity in latent spatial representations, contributing to advancing generative models in machine learning and artificial intelligence. The code that implements the proposal is available in https://github.com/marianorivera/How2TrainUrVAE.
@@ 2 Background
The goal is to learn the generative model of ${\mathcal{D}}$ a dataset with $N\geq 1$ independent and identically distributed (iid) data points: ${\mathcal{D}}=\{{\bf x}^{(i)}\}_{i=1}^{N}$ with ${\bf x}^{(i)}\in\mathbb{R}^{m}$; assuming a set of latent variables, one for each data point: ${\mathcal{Z}}=\{{\bf z}^{(i)}\}_{i=1}^{N}$ with ${\bf z}^{(i)}\in\mathbb{R}^{d}$, and $d<m$. In this way, we have $({\bf x},{\bf z})\sim p({\bf x},{\bf z})=p({\bf x}|{\bf z})p({\bf z})$, which suggests a two-step generative model:
@@
$\mathrm{ELBO}(\phi,\theta)=$ $\beta_{1}\,\mathrm{KL}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr{)}$ (3)
$-\beta_{2}\,\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]$ (4)
where the parameters $\beta_{1}$ and $\beta_{2}$ weigh the relative contribution of each term to the total cost. The standard VAE formulation uses $\beta_{1}=\beta_{2}=1$, meanwhile $\beta$-VAE propose to use $\beta_{1}>1$ and $\beta_{2}=1$ to improve interpretable factorised latent representations; i.e., to learn disentangled latents [9]. The second term (4) is the negative log-likelihood that enforces the generated data, $\hat{\bf x}^{(i)}$, to be likely to belong to the training dataset. The first term (3) is the Kullback-Leibler divergence that enforces the approximated posterior $q_{\phi}$ to be similar to a chosen prior $p_{\lambda}$; i.e. ${\bf z}\sim p_{\lambda}({\bf z})$ and $\lambda$ are the parameters of the prior. The most common approach is to assume the prior distribution $p_{\lambda}({\bf z})$ a multivariate Gaussian with a mean of zero and an identity covariance matrix:
@@
Hence, one can factorize the posterior as [10]:
@@
$=\prod_{{\bf x}^{(i)}\in\mathcal{D}}\prod_{j=1}^{d}\mathcal{N}({\bf z}_{j};\mu_{j}^{(i)},\sigma_{j}^{(i)}).$ (9)
This factorization takes advantage of the fact that the product of Gaussian (densities) is a Gaussian (density). The second row is the product of the individual posteriors (associated with each data point ${\bf x}^{i}$), and the third row makes explicit the marginals (of each latent dimension ${\bf x}_{j}$). Fig. 2 illustrates the Gaussian mixture model for the posterior. The dashed line represents the posterior $q_{\phi}$; i.e., the product of the individual Gaussian $q_{\phi}^{(i)}$ associated with each data point (in color). Note that while $q_{\phi}$ is enforced to be similar to the prior $p_{\lambda}$: a Gaussian with mean equal to zero and unitary variance. Since the KL between two univariate Gaussian is given by
@@
$\mathrm{KL}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr{)}=\sum_{j=1}^{d}\mathrm{KL}\bigl{(}\mathcal{N}(\bar{\mu}_{j},\bar{\sigma}_{j}^{2})\|\mathcal{N}(0,I)\bigr{)}\\ =\frac{1}{2}\sum_{j=1}^{d}\sum_{i:{\bf x}^{(i)}\in D}\left[(\sigma_{j}^{(i)})^{2}+(\mu^{(i)}_{j})^{2}-1-\log(\sigma_{j}^{(i)})^{2}\right];$ (11)
where $\bar{\mu}$ is the mean vector. The ELBO minimization is efficiently achieved by using stochastic gradient descent by using samples of the training set (mini-batches): $\mathcal{M}\subset\mathcal{D}$.
@@ 3.1 The posterior collapse problem and over–smoothed reconstructions
VAEs often suffer from the posterior collapse problem [11] (latent variable collapse or KL vanishing); such a problem occurs when $\mathrm{KL}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\lambda}(\mathbf{z})\bigr{)}\rightarrow 0$. The posterior collapse to the prior results in non-identifiable latent variables: the likelihood function is independent of the latent representation, producing a smoothness of the reconstructed images. To alleviate such problem there has been suggested many possible solutions, see Ref. [12] and references therein. Despite many efforts to alleviate the posterior collapse, one can yet observe over-smoothed predictions [5, 11, 12]. It seems that only hierarchical computational expensive VAE architectures can generate realistic textured data [13, 3].
To reduce the blurring, a naive strategy could be to reduce the parameter $\beta_{1}<1$ in (3) and pass the weight to the log-likelihood. This presents another problem that we will analyze following. Since the covariance $\bar{\sigma}$ of the posterior $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ is a diagonal (non–correlated) matrix, the product in (8) can be decomposed in products of univariate Gaussians, one product for each dimension $j$. Thus, the $j$–th product results in
@@
and $s_{j}$ is a scale factor whose exact form can be found in [14]. If to obtain reconstructions that are more faithful to the data, we weight more the log–likelihood term into the ELBO, then the individual posterior $q_{\phi}^{(i)}$ variance is reduced so the variations of the predictions are limited such that the VAE conducts close to a simple Autoencoder. In such case, Equations (12)–(14) make evident a problem that arises when the individual posterior $q_{\phi}^{(i)}$ approaches a Dirac delta: $\sigma_{j}^{(i)}\approx 0$. Thus, the $j$–th entry of the posterior $q_{\phi}$ [see (7)] approaches a Dirac Delta: $\bar{\sigma}_{j}\approx 0$. As a result, the $z_{j}$ latent variable is meaningless. Therefore, a significantly large $\beta_{1}$ value would be expected to result in a blurrier reconstruction.
@@ 3.2 The posterior as a mixture of Gaussians
To introduce our proposal, we first note that the latent variable ${\bf z}^{(i)}$, associated with the data point ${\bf x}^{(i)}$, is also a stochastic variable with its particular distribution, the individual posterior:
@@
where $q^{(i)}_{\phi}({\bf z})=\mathcal{N}(\mu^{(i)},\sigma^{(i)})$ with $\sigma^{(i)}\succ 0$. Such individual posteriors need to be combined to define the posterior. A direct possibility is factorization (7) but others factorization are possible. In this work, instead of product (7), we propose the mixture model:
@@
with $\alpha_{i}=1/N$ (with $N=\sharp\mathcal{D}$), which implies that all data points are equally relevant. Hence, the mean of the mixture is given by:
@@
Thus, we can estimate the mean and variance of the global posterior with a Stochastic Gradient strategy using data batches ${\cal M}\subset{\cal D}$.:
$\bar{\mu}_{j}$ $\approx\frac{1}{\sharp\mathcal{M}}\sum_{i:{\mathbf{x}}^{(i)}\in{\cal M}}\mu^{(i)}_{j},$ (18)
$\bar{\sigma}^{2}_{j}$ $\approx\frac{1}{\sharp\mathcal{M}}\sum_{i:{\mathbf{x}}^{(i)}\in{\cal M}}\left({\mathbf{z}}^{(i)}_{j}-\bar{\mu}_{j}\right)^{2}.$ (19)
@@ 3.3 Loss and training strategy
In this section, we present our loss function. First, we note that the classic $\mathrm{KL}$ formulation [1] encourages the individual means $\mu^{(i)}$’s to be zero, contributing to the so-called posterior collapse effect: the ${\mathbf{z}}^{(i)}$ are non–identifiable. Differently, in our proposal, we encourage the mean of those $\mu^{(i)}$’s to approach zero:
$\mathrm{KL}_{G}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})\,\|\,p_{\lambda}(\mathbf{z})\bigr{)}=\mathrm{KL}\bigl{(}\mathcal{N}(\bar{\mu},\bar{\Sigma})\,\|\,\mathcal{N}(0,I)\bigr{)}\\ =\frac{1}{2}\sum_{j=1}^{d}\left[\bar{\sigma}_{j}^{2}+\bar{\mu}_{j}^{2}-1-\log\bar{\sigma}_{j}^{2}\right];$ (20)
where we define covariance diagonal matrix of the approximated posterior as $\bar{\Sigma}=\mathrm{diag}(\bar{\sigma}^{2})$, and $\bar{\mu}$ and $\bar{\sigma}^{2}$ are given in (18) and (19), respectively.
Furthermore, one behavior we aim to prevent is the collapse of the variances of individual posteriors $(\sigma^{(i)})^{2}$ to zero. That eliminates the stochastic behavior of the Encoder, causing the VAE to behave like an AE. In such a scenario, the VAE would be limited to generalizing: to generate data not part of its training data [15]. Thus, we introduce a regularization term that penalizes individual variances that become close to zero, preventing such individual posteriors from becoming Dirac deltas. This regularization term is:
@@
Note that KL (21) does not penalize the individual means.
It is well known that metrics based on pixel-to-pixel comparisons are not the most appropriate for evaluating the quality of data produced by generating models. We will tolerate significant reconstruction errors as long as the reconstruction looks realistic. In the example we use, this means that we prefer to generate images where the hairstyle seems realistic, even if the positions of the curls in the reconstruction do not match the input image, rather than over-smoothing regions of the hair (as is common in a typical VAE). GAN-type models have been relatively successful in generating realistic images [16], despite suffering from the problem of modes collapse. Similar to VAE-GAN, which trains the VAE based on a GAN-type scheme [17], we also use a GAN training scheme. In particular, we use a PatchGAN strategy [18, 19, 20], which, instead of giving a reality measurement rating for the entire image, generates a two-dimensional array where each element evaluates a support region in the generated image or real.
Therefore, the likelihood term includes two metrics: a pixel-to-pixel measure that enforces the restoration to be similar to the input data and a general evaluation grade of the image realism. For the first term we choose the L1 norm known for being robust to errors in the exact reconstruction of structures’ positions in high-frequency texture areas. This is equivalent to said that the error follows a Laplace distribution. For the second term, we use a Convolutional Neural Network (Discriminator) that, given the restored image $\hat{\mathbf{x}}$, of size (H,W), produces an output ${\mathbf{r}}$, of size $(h/8,w/8)$, where each entry reflects the realism degree of its support region. The discriminator is represented by the conditional ${\mathbf{r}}\sim p_{\gamma}({\mathbf{r}}|{\mathbf{x}})$ where ${\mathbf{r}}\in[0,1]^{h/8\times w/8}$, and $\gamma$ are the parameters of the neural network. Then the log–likelihood takes the form
$\mathbb{E}_{q_{\lambda}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]=\beta_{3}\|{\mathbf{x}}-\hat{\mathbf{x}}\|-\beta_{4}\mathrm{H}\left[{\mathbf{1}},p_{\gamma}({\mathbf{r}}|\hat{\mathbf{x}})\right]$ (22)
@@
$\mathrm{ELBO}_{reg}(\phi,\theta)=\beta_{1}\,\mathrm{KL}_{G}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})\,\|\,p_{\lambda}(\mathbf{z})\bigr{)}\\ +\beta_{2}\,\mathrm{KL}_{I}\bigl{(}q_{\phi}(\mathbf{z}|\mathbf{x})\,\|\,p_{\lambda}(\mathbf{z})\bigr{)}\\ +\beta_{3}\|{\mathbf{x}}-\hat{\mathbf{x}}\|-\beta_{4}\mathrm{H}\left[{\mathbf{1}},p_{\gamma}({\mathbf{r}}|\hat{\mathbf{x}})\right]$ (23)
This loss encourages the generator to produce realistic predictions $\hat{\mathbf{x}}$. To train the discriminator, we minimize the loss
$L(\gamma)=-\beta_{4}\left(\mathrm{H}\left[{\mathbf{1}},p_{\gamma}({\mathbf{r}}|{\mathbf{x}})\right]+\mathrm{H}\left[{\mathbf{0}},p_{\gamma}({\mathbf{r}}|\hat{\mathbf{x}})\right]\right)$ (24)
Fig. 3: Residual-VAE components. (a) Residual encoder, (b) Gaussian sampler, and (c) residual decoder. In the presented experiments, we use an architecture with six residuals blocks ($R=6$) with two identity blocks each one ($r=2$).
In all the conducted experiments, we kept the weight parameters: $(\beta_{1},\beta_{2},\beta_{3},\beta_{4})=(1,0.5,5000,100)$. For computational efficiency purposes, we implement the training of both the VAE (generator) at the discriminator, using a stochastic descent strategy with batch size equal to $b=20$. Note that as the batch size $b$ becomes larger, the estimation of ELBO terms [Eqs. (20), (21) and (22)] becomes more precise. Therefore, avoiding using batches close to one is important to ensure stable convergence of the stochastic gradient. Algorithm 1 summarizes training details.
@@
2:     ${\mathbf{z}}\sim q_{\phi}({\mathbf{z}}|{\mathbf{x}})$triangleright Encode
3:     $\hat{\mathbf{x}}\sim p_{\theta}({\mathbf{x}}|{\mathbf{z}})$triangleright Decode
4:     ${\mathbf{r}}\sim p_{\gamma}({\mathbf{r}}|\hat{\mathbf{x}})$triangleright Realism scores
@@
6:     Take a gradient descent step on $\nabla_{\gamma}L(\gamma)$.